OpenStudy (anonymous):

The displacement from equilibrium of a mass oscillating on the end of a spring suspended from a ceiling is given below, where y is the displacement (in feet) and t is the time in seconds. Use a graphing utility to graph the displacement function on the interval [0, 10]. Find the value of t past which the displacement is less than 3 inches from equilibrium. (Round your answer to two decimal places.) y = [1.54e^(−0.22t)]*cos(4.9t)

OpenStudy (whpalmer4):

\[y = 1.54e^{-0.22t}\cos(4.9t)\] You want the value of \(t\) where \(y < \frac{3}{12}\) \[\frac{3}{12} > 1.54e^{-0.22t}\cos(4.9t)\] Key here is to realize that \(\cos(4.9t)\) oscillates between 1 and -1. We can replace that part with 1 to find the point at which the damping exponential function is a small enough value that we'll never be more than 3 inches from equilibrium.

OpenStudy (whpalmer4):

Here's a graph of a similar equation (numbers changed, so don't just copy and paste this into your answer!) Two gridlines show the amplitude of 3 inches.

OpenStudy (anonymous):

Where do you get\[\frac{ 3 }{ 12 }\]from?

OpenStudy (whpalmer4):

" y is the displacement (in feet)" "displacement is less than 3 inches from equilibrium"

OpenStudy (whpalmer4):

3 inches = 3/12 of a foot, right?

OpenStudy (anonymous):

So then what does t equal?

OpenStudy (anonymous):

t=nπ4.9

OpenStudy (anonymous):

t=nPi/4.9

OpenStudy (anonymous):

oops lol

OpenStudy (anonymous):

What is n then?

OpenStudy (anonymous):

The answer is suppose to be a number

OpenStudy (whpalmer4):

Yes. You need to find the value of \(t\) so that the "envelope" function (the exponential) is going to be less than the cutoff value for the displacement. Remember, the cos function ranges between -1 and 1.

OpenStudy (whpalmer4):

The dashed line is the exponential "envelope" function that is multiplied by the cos function. When the envelope gets down to 3/12, no matter what the value of the cos function is, the resulting displacement will be less than 3 inches.

OpenStudy (anonymous):

So do I put \[\int\limits_{0}^{3}1.54e ^{-.22t}\cos(4.9t) dt\]?

OpenStudy (whpalmer4):

No...Solve the following: \[\frac{3}{12} = 1.54e^{-0.22t}\]That's all there is to it...

OpenStudy (whpalmer4):

All you are trying to do is find the value of \(t\) where the dashed line crosses the horizontal line in this figure:

OpenStudy (anonymous):

I got 8.26 but still got it wrong

OpenStudy (whpalmer4):

If you've copied the problem correctly, that should be correct. Beyond that value of \(t\), it is impossible for the displacement to be greater than 3 inches.

OpenStudy (anonymous):

I already put it on my calculator: 1.54e^(−0.22*8.26399) = .25 = 3/12

OpenStudy (anonymous):

Why does Cos(4.9t) get left out?

OpenStudy (whpalmer4):

Because it can't be larger than 1. At its maximum value, the displacement is just the value of that exponential.